### v1

Abstract
Quasi-uniformity is a fundamental geometric property of point sets, crucial for applications such as kernel interpolation, Gaussian process regression, and space-filling experimental designs. While quasi-Monte Carlo methods are widely recognized for their low-discrepancy characteristics, understanding their quasi-uniformity remains important for practical applications. For the two-dimensional Sobol’ sequence, Sobol’ and Shukhman (2007) conjectured that the separation radius of the first $N$ points achieves the optimal rate $N^{-1/2}$, which would imply quasi-uniformity. This conjecture was disproved by Goda (2024), who computed exact values of the $\ell^{2}$-separation radius for a sparse subsequence $N=2^{2^{v}-1}$. However, the general behavior of the Sobol’ sequence for arbitrary $N$ remained unclear. In this paper, we derive exact expressions for the $\ell^{\infty}$-separation radius of the first $N=2^{m}$ points of the two-dimensional Sobol’ sequence for all $m\in\mathbb{N}$. As an immediate consequence, we show that the separation radius of Sobol’ points is $O(N^{-3/4})$, which is strictly worse than the optimal rate $N^{-1/2}$, revealing that the two-dimensional Sobol’ sequence has a suboptimal mesh ratio that grows at least as $N^{1/4}$.
Keywords: Quasi-Monte Carlo, Sobol’ sequence, space-filling design, covering radius, separation radius, mesh ratio
AMS subject classifications: 05B40, 11K36, 52C15
1 Introduction
Quasi-Monte Carlo (QMC) methods replace random sampling with carefully constructed deterministic point sets for numerical integration over the unit cube; see, e.g., [4, 5, 6, 13, 14, 19]. While classical QMC theory focuses on discrepancy, which measures deviation from perfect equidistribution, it does not directly control local spacing between points. However, many applications require a more geometric notion of uniformity. In particular, tasks such as scattered data approximation [23], Gaussian process regression [22], kernel interpolation [18], and the design of computer experiments [7, 15, 17] benefit from point sets that ensure both well-controlled spacing and coverage. These properties are summarized by the term quasi-uniformity.
To formalize quasi-uniformity, two standard geometric parameters are commonly used for a finite point set $Q\subset[0,1]^{d}$: the covering radius $h_{p}(Q)$ and the separation radius $q_{p}(Q)$. Specifically, for the $\ell^{p}$ norm $\|\cdot\|_{p}$, define
$h_{p}(Q)\ :=\sup_{x\in[0,1]^{d}}\min_{y\in Q}\|x-y\|_{p},\qquad q_{p}(Q)\ :=\dfrac{1}{2}\min_{\begin{subarray}{c}x,y\in Q\\ x\neq y\end{subarray}}\|x-y\|_{p}.$
Intuitively, $h_{p}(Q)$ is the smallest radius such that closed $\ell^{p}$ balls around each point cover $0,1]^{d}$, whereas $q_{p}(Q)$ is the largest radius such that the corresponding open balls around the points do not overlap. As noted in [14, Section 6], [21], and [16], this geometric interpretation implies
$h_{p}(Q)\in\Omega(|Q|^{-1/d}),\qquad q_{p}(Q)\in O(|Q|^{-1/d}).$ (1)
The mesh ratio
$\rho_{p}(Q):=\dfrac{h_{p}(Q)}{q_{p}(Q)}$
quantifies how close $Q$ is to an ideal packing/covering configuration. A sequence $(\boldsymbol{x}_{n})_{n\geq 0}$ is quasi-uniform in $\ell_{p}$ if the mesh ratio $\rho_{p}(Q_{N})$, for $Q_{N}=\{\boldsymbol{x}_{0},\dots,\boldsymbol{x}_{N-1}\}$, is bounded independently of $N$. Equivalently, by (1), both $h_{p}(Q_{N})$ and $q_{p}(Q_{N})$ are $\Theta(N^{-1/d})$.
Given that QMC point sets are natural candidates for generating uniform point sets, it is important to ask whether they are quasi-uniform. As noted in [24], however, the quasi-uniformity of classical QMC constructions had long been unresolved. This changed with Goda’s influential result [9], which shows that the two-dimensional Sobol’ sequence is not quasi-uniform.
The Sobol’ sequence [20] is widely used in practice due to its efficient digital construction, extensibility to arbitrary sample sizes, and effectiveness in high-dimensional integration. It is available in standard software packages, such as Python’s QMCPy [1] and MATLAB’s Statistics and Machine Learning Toolbox. Sobol’ and Shukhman [21] conjectured that the separation radius of the first $N$ points of the $d$-dimensional Sobol’ sequence behaves like $N^{-1/d}$, which would imply quasi-uniformity. For $d=2$, Goda [9] showed that this conjecture fails for the sparse subsequence of the form $N=2^{2^{w}-1}$, and the case $N=2^{2^{w}}$ was subsequently analyzed in [3]. Nevertheless, a complete description of the separation radius for general $N$ remains open, and addressing this gap is the focus of the present work.
Main results. Let $Q_{N}$ denote the first $N$ points of the two-dimensional Sobol’ sequence (Definition 2.1). Our main contributions are as follows:
  • We provide an exact formula for the $\ell^{\infty}$-separation radius of the dyadic prefixes $Q_{2^{m}}$.
  • We prove that $q_{\infty}(Q_{N})\in O(N^{-3/4})$ and $\rho_{\infty}(Q_{N})\in\Omega(N^{1/4})$.
These results are summarized in the following theorem and its corollaries.
Theorem 1.1.
Let $m\in\mathbb{N}$ and $N=2^{m}$. Let $Q_{N}$ be the first $N$ points of the two-dimensional Sobol’ sequence. If $m=2^{v}$ or $2^{v}-1$ for some $v\in\mathbb{N}$, then we have
$q_{\infty}(Q_{N})=2^{-m-1}.$
Otherwise, decomposing $m$ as $m=2^{v}+2^{w}+c$ with integers $v>w\geq 0$ and $0\leq c<2^{w}$, then
$q_{\infty}(Q_{N})=2^{-{2^{v}}-2^{w}}.$
The following corollaries show that the separation radius of the two-dimensional Sobol’ sequence decays as $O(N^{-3/4})$, which is a factor $N^{-1/4}$ smaller than the optimal order $\Theta(N^{-1/2})$. As a consequence, the Sobol’ sequence has a suboptimal mesh ratio for all $N$. Since all $\ell^{p}$ norms on $\mathbb{R}^{2}$ are equivalent, the asymptotic orders stated below remain valid for any $p\in[1,\infty]$.
Corollary 1.2.
For any $m\in\mathbb{N}$, we have
$q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-5/4},$ (2)
$q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-3/2}\quad\text{for }m\neq 1,5,$ (3)
with equality in the second inequality if $m=2^{v}-2$ for some $v\geq 2$.
Corollary 1.3.
For any integer $N\geq 2$, the following bounds hold:
$q_{\infty}(Q_{N})$ $\leq C_{1}N^{-3/4},$ (4)
$\rho_{\infty}(Q_{N})$ $\geq C_{2}N^{1/4},$ (5)
where $C_{1}=C_{2}=2^{-1/2}$. Moreover, for $N\geq 64$, the constants can be improved to $C_{1}=2^{-3/4}$ and $C_{2}=2^{-1/4}$.
Related work. In recent years, the quasi-uniformity of QMC point sets has been the subject of intensive investigation. There are two major classes of QMC constructions: digital nets and sequences [6, 14], and lattice point sets (including their infinite analogue, Kronecker sequences) [4, 19].
For lattice point sets, quasi-uniformity has been actively studied. In one dimension, the separation radius of Kronecker sequences is completely characterized in [8]. More general results in higher dimensions are provided in [2], including bounded mesh ratios for two-dimensional Fibonacci lattices, existence results for $d$-dimensional lattice rules, and explicit constructions for $d$-dimensional Kronecker sequences.
Turning to digital nets and sequences, the one-dimensional van der Corput sequence in base $b$ is known to be quasi-uniform since its first $b^{m}$ points are $\{i/b^{m}\mid 0\leq i<b^{m}\}$. In higher dimensions, the covering radius, often referred to as dispersion, has been extensively studied [14, Chapter 6]. In particular, for $(t,d)$-sequences in base bb—well-known examples include the Sobol’, Faure, and Niederreiter sequences—, the covering radius is known to attain the optimal order $\Theta(N^{-1/d})$ for any dimension $d$. Thus, the problem of establishing quasi-uniformity reduces to verifying whether the separation radius also scales as $N^{-1/d}$.
For $d=2$, as stated, the Sobol’ sequence is not quasi-uniform [9]. The separation radius of several two-dimensional digital nets was studied in [10, 11]. Numerical experiments therein suggest that the Larcher–Pillichshammer nets [12] are quasi-uniform. This was theoretically proved by Dick, Goda, and Suzuki [3], who introduced an algebraic criterion for well-separated digital nets. To our knowledge, this remains the only explicit construction of low-discrepancy and quasi-uniform digital nets for $d\geq 2$. The paper [3] also shows the non-optimality of the separation radius for some two-dimensional digital nets and Fibonacci polynomial lattices, as well as for $b$-dimensional Faure sequences in prime base $b$.
From a different perspective, Pronzato and Zhigljavsky [16] constructed quasi-uniform infinite sequences via a greedy packing algorithm, ensuring a mesh ratio of at most $2$ for the first $N$ points, $N\geq 2$. However, these sequences do not necessarily maintain low discrepancy in dimensions $d\geq 2$.
Organization. Preliminaries and notation are collected in Section 2. Section 3 provides the necessary lemmas, the proof of Theorem 1.1, and the derivation of Corollaries 1.2 and 1.3.
2 Preliminaries
Notation. Throughout this paper, let $\mathbb{F}_{2}$ denote the finite field of order $2$, $\mathbb{N}$ the set of positive integers, and $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. Addition in $\mathbb{F}_{2}$ or $\mathbb{F}_{2}^{m}$ is denoted by $\oplus$. We write $\boldsymbol{0}_{m}\in\mathbb{F}_{2}^{m}$ for the zero vector and $\boldsymbol{1}_{m}\in\mathbb{F}_{2}^{m}$ for the all-ones vector. The subscript $m$ will be omitted whenever it does not cause confusion.
For an integer $0\leq n<2^{m}$ with binary expansion
$n=n_{1}+2n_{2}+\cdots+2^{m-1}n_{m},$
we define
$\vec{n}=(n_{1},n_{2},\ldots,n_{m})^{\top}\in\mathbb{F}_{2}^{m},$
where $n_{1},\dots,n_{m}\in\{0,1\}$ are identified with elements of $\mathbb{F}_{2}$.
For a vector $\boldsymbol{z}=(z_{1},\dots,z_{m})^{\top}\in\mathbb{F}_{2}^{m}$, we denote by $\boldsymbol{z}[i]\in\{0,1\}$ its $i$th component $z_{i}$, and for $i<j$ we define the slice $\boldsymbol{z}[i\!:\!j]=(z_{i},\dots,z_{j})^{\top}$. Finally, we set
$\phi(\boldsymbol{z})=\boldsymbol{z}[1]2^{-1}+\cdots+\boldsymbol{z}[m]2^{-m}.$
For a matrix $P=(P_{ij})_{1\leq i,j\leq m}$, we use the notation $P[i][j]:=P_{ij}$ for convenience, and define
$P[x\!:\!y][z\!:\!w]:=(P[i][j])_{x\leq i\leq y,\;z\leq j\leq w},$
which represents the submatrix of $P$ consisting of rows $x$ through $y$ and columns $z$ through $w$.
Pascal matrix. The (upper triangular) Pascal matrix $P_{m}\in\mathbb{F}_{2}^{m\times m}$ is defined by
$P_{m}[i][j]\equiv\binom{j-1}{i-1}\pmod{2},\quad 1\leq i,j\leq m.$ (6)
The subscript $m$ will be omitted whenever it does not cause confusion.
Two-dimensional Sobol’ sequence. The two-dimensional Sobol’ sequence is defined as follows. For the definition in general dimension $d$, we refer the reader to [6, Chapter 8].
Definition 2.1.
Let $n\in\mathbb{N}_{0}$ and choose $m\in\mathbb{N}$ such that $n<2^{m}$. The two-dimensional Sobol’ sequence $(\boldsymbol{x}_{n})_{n\in\mathbb{N}}$ is a sequence of points in $0,1)^{2}$, where the $n$th point is given by
$\boldsymbol{x}_{n}:=(\phi(\vec{n}),\phi(P_{m}\vec{n})).$
This definition does not depend on the choice of $m$: increasing $m$ simply pads $\vec{n}$ with leading zeros, and since $P_{m}$ is upper triangular, $\phi(\vec{n})$ and $\phi(P_{m}\vec{n})$ remain unchanged.
This construction is generalized to the notions of digital nets and digital sequences. The uniformity of these point sets is usually measured by a quantity called the $t$-value (see [14, Chapter 4] or [6, Chapter 4]). It is known that the $t$-value of the two-dimensional Sobol’ sequence attains the best possible value, namely zero [6, Section 8.1]. This discussion can be formulated rigorously in the following proposition.
Proposition 2.2.
Let $m\in\mathbb{N}$ and let $Q_{2^{m}}$ denote the set of the first $2^{m}$ points of the two-dimensional Sobol’ sequence. Then, for any rectangle of the form
$\left[\frac{a}{2^{k}},\frac{a+1}{2^{k}}\right)\times\left[\frac{b}{2^{l}},\frac{b+1}{2^{l}}\right),\quad(a,b,k,l\in\mathbb{N}_{0},\,0\leq a<2^{k},\,0\leq b<2^{l})$
with $k+l=m$, there is exactly one point from $Q_{2^{m}}$ contained in the rectangle.
3 Proofs
3.1 Lemmas
We make heavy use of the properties of the Pascal matrix. In particular, its entries modulo $2$ can be characterized using Lucas’s theorem. Specifically, for integers $0\leq p,q<2^{m}$, we have
$\binom{p}{q}\equiv\prod_{i=1}^{m}\binom{\vec{p}[i]}{\vec{q}[i]}\pmod{2}$. (7)
Using this result, we can establish the following properties.
Lemma 3.1.
Let $P$ be the Pascal matrix. Let $v>w\geq 0$ and $i\geq 0$ be integers, and set $V:=2^{v}$ and $W:=2^{w}$. Then the following hold:
  1. (i)
    $P[i][W]=1\iff 1\leq i\leq W$.
  2. (ii)
    $P[i][V+W]=1\iff 1\leq i\leq W\text{ or }V+1\leq i\leq V+W$.
  3. (iii)
    For any $\vec{p}\in\mathbb{F}_{2}^{m}$ with $m\leq 2V-1$, we have $(P_{m}\vec{p})[V]=\vec{p}[V]$ and $(P_{m}\vec{p})[V+W]=\vec{p}[V+W]$.
  4. (iv)
    For any $\vec{p}\in\mathbb{F}_{2}^{m}$ with $m\leq 2V-2$, we have $(P_{m}\vec{p})[V-1]=\vec{p}[V-1]\oplus\vec{p}[V]$.
  5. (v)
    $P_{V}=P[1\!:\!V][V+1\!:\!2V]=P[V+1\!:\!2V][V+1\!:\!2V]$.
  6. (vi)
    $(P_{W}\boldsymbol{1}_{W})[i]=1\iff i=W$.
  7. (vii)
    $(P_{V+W}\boldsymbol{1}_{V+W})[i]=1\iff i=W,V\text{ or }V+W$.
Proof.
Items (i)(v) follow directly from Lucas’s theorem (7).
To prove the remaining items, we note that for any $m\in\mathbb{N}$ and $1\leq i\leq m$, the hockey-stick identity implies
$(P_{m}\boldsymbol{1}_{m})[i]\equiv\bigoplus_{j=1}^{m}\binom{j-1}{i-1}\equiv\binom{m}{i}\pmod{2}$.
Using this fact, Items (vi) and (vii) also follow from Lucas’s theorem. ∎
The following results assert that the binary representations of two close points are related.
Lemma 3.2.
Let $\ell,m,p,q$ be integers with $2\leq\ell\leq m$ and $0\leq p,q<2^{m}$. Assume that $0\leq\phi(\vec{q})-\phi(\vec{p})<2^{-\ell+1}$. Then one of the following holds:
  1. (i)
    $\vec{p}[1\!:\!\ell-1]=\vec{q}[1\!:\!\ell-1]$.
  2. (ii)
    There exists $1\leq k\leq\ell-1$ such that all of the following conditions hold:
    1. (a)
      $\vec{p}[1\!:\!k-1]=\vec{q}[1\!:\!k-1]$,
    2. (b)
      $\vec{p}[k]=0$, $\vec{q}[k]=1$,
    3. (c)
      $\vec{p}[k+1\!:\!\ell-1]=\boldsymbol{1}$, $\vec{q}[k+1\!:\!\ell-1]=\boldsymbol{0}$,
    4. (d)
      $\vec{p}[\ell]\geq\vec{q}[\ell]$,
    5. (e)
      $\vec{p}[\ell\!:\!m]\neq\vec{q}[\ell\!:\!m]$.
Proof.
We assume that $\vec{p}[1\!:\!\ell-1]\neq\vec{q}[1\!:\!\ell-1]$, since otherwise there is nothing to prove. Let $k$ be the smallest index with $1\leq k\leq\ell-1$ such that $\vec{p}[k]\neq\vec{q}[k]$. Since we have assumed that $\phi(\vec{p})<\phi(\vec{q})$, it follows that $\vec{p}[k]=0$ and $\vec{q}[k]=1$. By the minimality of $k$, we also have $\vec{p}[1\!:\!k-1]=\vec{q}[1\!:\!k-1]$. Thus, (ii)(a) and (ii)(b) are established.
We now prove (ii)(c) by contradiction. Assume that there exists $k<k^{\prime}\leq\ell-1$ such that $\vec{p}[k^{\prime}]=0$ or $\vec{q}[k^{\prime}]=1$. Then, using (ii)(a) and (ii)(b), we have
$\phi(\vec{q})-\phi(\vec{p})=2^{-k}+\sum_{i=k+1}^{m}(\vec{q}[i]-\vec{p}[i])2^{-i}\geq 2^{-k}+2^{-k^{\prime}}-\sum_{i=k+1}^{m}2^{-i}=2^{-k^{\prime}}+2^{-m},$
which contradicts the assumption that $\phi(\vec{q})-\phi(\vec{p})<2^{-\ell+1}$.
The proofs of (ii)(d) and (ii)(e) are similar to that of (ii)(c) and are omitted. ∎
In particular, this lemma implies the following corollary.
Corollary 3.3.
Let $\ell,m,p,q$ be integers with $2\leq\ell\leq m$ and $0\leq p\neq q<2^{m}$, and assume that $|\phi(\vec{q})-\phi(\vec{p})|<2^{-\ell+1}$. Then the following statements hold:
  1. (i)
    The vector $(\vec{p}\oplus\vec{q})[1\!:\!\ell-1]$ is either $\boldsymbol{0}$, $\boldsymbol{1}$, or of the form $(0,\dots,0,1,\dots,1)^{\top}$.
  2. (ii)
    If there exists an integer $2\leq k\leq\ell-1$ such that $(\vec{p}\oplus\vec{q})[k-1]=0$ and $(\vec{p}\oplus\vec{q})[k]=1$, then
    $(\vec{p}\oplus\vec{q})[1\!:\!k-1]=\boldsymbol{0},\quad(\vec{p}\oplus\vec{q})[k\!:\!\ell-1]=\boldsymbol{1},\quad(\vec{p}\oplus\vec{q})[\ell\!:\!m]\neq\boldsymbol{0}.$
    Moreover, one of the following holds:
    • $\phi(\vec{p})>\phi(\vec{q})$, $\vec{p}[k]=1$, $\vec{q}[k]=0$, $\vec{p}[k+1\!:\!\ell-1]=\boldsymbol{0}$, $\vec{q}[k+1\!:\!\ell-1]=\boldsymbol{1}$, $\vec{p}[\ell]\leq\vec{q}[\ell]$;
    • $\phi(\vec{p})<\phi(\vec{q})$, $\vec{p}[k]=0$, $\vec{q}[k]=1$, $\vec{p}[k+1\!:\!\ell-1]=\boldsymbol{1}$, $\vec{q}[k+1\!:\!\ell-1]=\boldsymbol{0}$, $\vec{p}[\ell]\geq\vec{q}[\ell]$.
  3. (iii)
    If there exists an integer $2\leq k\leq\ell-1$ such that $(\vec{p}\oplus\vec{q})[k-1]=(\vec{p}\oplus\vec{q})[k]=1$, then
    $(\vec{p}\oplus\vec{q})[k\!:\!\ell-1]=\boldsymbol{1},\quad(\vec{p}\oplus\vec{q})[\ell\!:\!m]\neq\boldsymbol{0}.$
    Moreover, one of the following holds:
    • $\phi(\vec{p})>\phi(\vec{q})$, $\vec{p}[k\!:\!\ell-1]=\boldsymbol{0}$, $\vec{q}[k\!:\!\ell-1]=\boldsymbol{1}$, $\vec{p}[\ell]\leq\vec{q}[\ell]$;
    • $\phi(\vec{p})<\phi(\vec{q})$, $\vec{p}[k\!:\!\ell-1]=\boldsymbol{1}$, $\vec{q}[k\!:\!\ell-1]=\boldsymbol{0}$, $\vec{p}[\ell]\geq\vec{q}[\ell]$.
3.2 Proof of the main theorem
  • The case $m=2^{v}-1$ is essentially treated in [9]. As shown in the proof of [9, Theorem 2.2], we have
    $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}-1}\|_{\infty}=2^{-m}.$
    On the other hand, by the construction of the Sobol’ sequence, $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}\geq 2^{-m}$ holds for any $p\neq q$. Combining these observations, we conclude that
    $q_{\infty}(Q_{2^{m}})=2^{-m-1}.$
  • The case $m=2^{v}$ is essentially treated in [3]. From the proof of [3, Theorem 4.2] with $L$ being the identity matrix, we deduce $\boldsymbol{x}_{2^{m}+1}=(1/2+1/2^{m},1/2+1/2^{m})$, which implies $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}+1}\|_{\infty}=2^{-m}$. Hence, in the same manner as the previous case, we obtain
    $q_{\infty}(Q_{2^{m}})=2^{-m-1}.$
  • For the remaining cases, we exclude $m=2^{v}$ or $m=2^{v}-1$ for any $v\in\mathbb{N}$. We may then write
    $m=2^{v}+2^{w}+c,\quad v\geq 2,\quad v>w,\quad 2^{w}>c\geq 0,$
    and set $V=2^{v}$, $W=2^{w}$. Since $m+1$ is not a power of two, we have $m\leq 2V-2$ and can thus apply Lemma 3.1 (iii) and (iv). In the following, we will establish separately that
    $q_{\infty}(Q_{2^{m}})\leq 2^{-V-W}\quad\text{and}\quad 2^{-V-W}\leq q_{\infty}(Q_{2^{m}}).$
3.2.1 Proof of the upper bound
Let $p=2^{V+W-1}+2^{W-1}$ and $q=2^{V+W}-2^{W}$. To prove $q_{\infty}(Q_{2^{m}})\leq 2^{-V-W}$, it suffices to show that $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-V-W+1}$.
First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=W$ or $i=V+W$, it follows from Lemma 3.1 (i) and (ii) that
$x_{p,1}$ $=\phi(\vec{p})=2^{-W}+2^{-V-W},$
$x_{p,2}$ $=\phi(P_{m}\vec{p})=\sum_{i=1}^{V+W}(P[i][W]\oplus P[i][V+W])2^{-i}=\sum_{i=V+1}^{V+W}2^{-i}=2^{-V}-2^{-V-W}.$
Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ for $W+1\leq i\leq V+W$, we have
$P_{m}\vec{q}$ $=P_{m}(\mathrm{concat}(\boldsymbol{1}_{W},\boldsymbol{0}_{m-W})\oplus\mathrm{concat}(\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}))$
$=\mathrm{concat}(P_{W}\boldsymbol{1}_{W},\boldsymbol{0}_{m-W})\oplus\mathrm{concat}(P_{V+W}\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}),$
where “concat” denotes vertical concatenation of column vectors. Thus, by Lemma 3.1 (vi) and (vii), we obtain
$x_{q,1}$ $=\phi(\vec{q})=\sum_{i=W+1}^{V+W}2^{-i}=2^{-W}-2^{-V-W},$
$x_{q,2}$ $=\phi(P_{m}\vec{q})=2^{-V}+2^{-V-W}.$
Hence,
$|x_{p,1}-x_{q,1}|=|x_{p,2}-x_{q,2}|=2^{-V-W+1},$
which gives the desired result. ∎
3.2.2 Proof of the lower bound
We prove the bound by contradiction. Assume that there exist integers $p,q$ with $0\leq p\neq q<2^{m}$ such that $\phi(\vec{p})<\phi(\vec{q})$ and $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}<2^{-V-W+1}$. Let $\vec{\Delta}:=\vec{p}\oplus\vec{q}$. We divide the analysis into three cases according to the values of $\vec{\Delta}[V]$ and $\vec{\Delta}[V-1]$.
Case 1: $\vec{\Delta}[V]=0$.
In this case, Lemma 3.1 (iii) gives
$(P\vec{\Delta})[V]=\vec{\Delta}[V]=0.$
Then, Corollary 3.3 (i) implies
$\vec{\Delta}[1\!:\!V]=(P\vec{\Delta})[1\!:\!V]=\boldsymbol{0}.$
Hence, $\boldsymbol{x}_{p}$ and $\boldsymbol{x}_{q}$ lie in the same interval of the form $a/2^{V},(a+1)/2^{V})\times[b/2^{V},(b+1)/2^{V})$ for some $a,b$ with $0\leq a,b<2^{V}$. By Proposition 2.2, this forces $\boldsymbol{x}_{p}=\boldsymbol{x}_{q}$, contradicting the assumption $p\neq q$.
Case 2: $\vec{\Delta}[V]=1$ and $\vec{\Delta}[V-1]=0$.
In this case, Corollary 3.3 (ii), together with the assumption $\phi(\vec{p})<\phi(\vec{q})$, implies
$\vec{\Delta}[V\!:\!V+W-1]=\boldsymbol{1},$ (8)
$\vec{\Delta}[V+W\!:\!m]\neq\boldsymbol{0},$ (9)
$\vec{p}[V]=0,\quad\vec{q}[V]=1,$ (10)
$\vec{p}[V+W]\geq\vec{q}[V+W].$ (11)
By Lemma 3.1 (iii) and (10), we have
$(P\vec{p})[V]=\vec{p}[V]=0,\quad(P\vec{q})[V]=\vec{q}[V]=1,\quad(P\vec{\Delta})[V]=\vec{\Delta}[V]=1,$
and from Lemma 3.1 (iv),
$(P\vec{\Delta})[V-1]=\vec{\Delta}[V-1]\oplus\vec{\Delta}[V]=1.$
Hence, since $(P\vec{p})[V]=0$ holds, the first alternative of Corollary 3.3 (iii) gives
$(P\vec{\Delta})[V\!:\!V+W-1]$ $=\boldsymbol{1},$ (12)
$(P\vec{p})[V+W]$ $\leq(P\vec{q})[V+W].$ (13)
We further divide the analysis into the following two subcases.
Case 2-1: $\vec{\Delta}[V+W]=0$.
In this case, using (8) and Lemma 3.1 (v), we have
$(P\vec{\Delta})[V+1\!:\!V+W-1]$ $=P[V+1\!:\!V+W-1][V+1\!:\!V+W-1]\cdot\vec{\Delta}[V+1\!:\!V+W-1]$
$\quad\oplus P[V+1\!:\!V+W-1][V+W\!:\!V+W]\cdot\vec{\Delta}[V+W]$
$\quad\oplus P[V+1\!:\!V+W-1][V+W+1\!:\!m]\cdot\vec{\Delta}[V+W+1\!:\!m]$
$=P_{W-1}\boldsymbol{1}\oplus\boldsymbol{0}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m]$
$=\boldsymbol{1}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m],$
where $P^{\prime}:=P[V+1\!:\!V+W-1][V+W+1\!:\!m]$. Combined with (12), this gives
$P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m]=\boldsymbol{0}.$ (14)
From Lemma 3.1 (v), we have
$P^{\prime}=P[V+1\!:\!V+W-1][V+W+1\!:\!m]$ $=P[1\!:\!W-1][W+1\!:\!m-V]$
$=P[1\!:\!W-1][1\!:\!m-V-W].$
Since $P_{m-V-W}$ is non-singular and $W-1\geq m-V-W$, the columns of $P^{\prime}$ are linearly independent. Hence, (14) implies
$\vec{\Delta}[V+W+1\!:\!m]=\boldsymbol{0},$
which together with the assumption $\vec{\Delta}[V+W]=0$ contradicts (9).
Case 2-2: $\vec{\Delta}[V+W]=1$.
In this case, (11) implies that $\vec{p}[V+W]=1$ and $\vec{q}[V+W]=0$. Hence, Lemma 3.1 (iii) implies
$(P\vec{p})[V+W]=\vec{p}[V+W]=1,\quad\text{and}\quad(P\vec{q})[V+W]=\vec{q}[V+W]=0,$
which contradicts (13).
Case 3: $\vec{\Delta}[V]=1$ and $\vec{\Delta}[V-1]=1$.
In this case, Corollary 3.3 (iii), combined with the assumption $\phi(\vec{p})<\phi(\vec{q})$, implies
$\vec{\Delta}[V\!:\!V+W-1]=\boldsymbol{1},$ (15)
$\vec{\Delta}[V+W\!:\!m]\neq\boldsymbol{0},$ (16)
$\vec{p}[V]=1,\vec{q}[V]=0,$ (17)
$\vec{p}[V+W]\geq\vec{q}[V+W].$ (18)
By Lemma 3.1 (iii) and (17), we have
$(P\vec{p})[V]=\vec{p}[V]=1,\quad(P\vec{q})[V]=\vec{q}[V]=0,\quad(P\vec{\Delta})[V]=\vec{\Delta}[V]=1,$
and from Lemma 3.1 (iv),
$(P\vec{\Delta})[V-1]=\vec{\Delta}[V-1]\oplus\vec{\Delta}[V]=0.$
Hence, since $(P\vec{p})[V]=1$ holds, the first alternative of Corollary 3.3 (ii) gives
$(P\vec{\Delta})[V\!:\!V+W-1]$ $=\boldsymbol{1},$ (19)
$(P\vec{p})[V+W]$ $\leq(P\vec{q})[V+W].$ (20)
We now split the analysis into the following two subcases.
Case 3-1: $\vec{\Delta}[V+W]=0$.
In this case, in the same way as in the proof of Case 2-1, (15) and (19) imply
$\vec{\Delta}[V+W+1\!:\!m]=\boldsymbol{0}.$
This, together with the assumption $\vec{\Delta}[V+W]=0$, contradicts (16).
Case 3-2: $\vec{\Delta}[V+W]=1$.
Here, (18) implies $\vec{p}[V+W]=1$ and $\vec{q}[V+W]=0$. Hence, Lemma 3.1 (iii) implies
$(P\vec{p})[V+W]=\vec{p}[V+W]=1,\quad\text{and}\quad(P\vec{q})[V+W]=\vec{q}[V+W]=0,$
which contradicts (20).
The proof is therefore complete in all cases. ∎
3.3 Proof of Corollary 1.2
The cases $m=1$ and $m=5$ hold individually, as in Theorem 1.1.
If $m=2^{v}$ or $2^{v}-1$ for some $v\in\mathbb{N}$ and $m\neq 1$, then Theorem 1.1 gives $q_{\infty}(Q_{2^{m}})=2^{-m-1}$, and hence
$2^{3m/4}q_{\infty}(Q_{2^{m}})=2^{-m/4-1}\leq 2^{-3/2}.$
Otherwise, write $m=2^{v}+2^{w}+c$ with $v>w$ and $2^{w}>c\geq 0$. Then Theorem 1.1 gives $q_{\infty}(Q_{2^{m}})=2^{-2^{v}-2^{w}}$.
First, consider $w\leq v-2$. Since $m\neq 5$, we have $v\geq 3$. Using $c\leq 2^{w}-1$, we obtain
$\log_{2}(2^{3m/4}q_{\infty}(Q_{2^{m}}))=\frac{3}{4}(2^{v}+2^{w}+c)-2^{v}-2^{w}\leq-\frac{2^{v}}{4}+\frac{2^{w}}{2}-\frac{3}{4}.$
Further, using $w\leq v-2$, we have $2^{w}/2\leq 2^{v-3}=2^{v}/8$, so that
$\log_{2}(2^{3m/4}q_{\infty}(Q_{2^{m}}))\leq-\frac{2^{v}}{4}+\frac{2^{v}}{8}-\frac{3}{4}=-\frac{2^{v}}{8}-\frac{3}{4}<-\frac{3}{2}.$
Next, consider w=v1w=v-1. In this case, $c\leq 2^{w}-2$; otherwise $m+1$ would be a power of two. Then
$\log_{2}(2^{3m/4}q_{\infty}(Q_{2^{m}}))$ $=\frac{3}{4}(2^{v}+2^{w}+c)-2^{v}-2^{w}$
$\leq\frac{3}{4}(2^{w+1}+2^{w}+2^{w}-2)-2^{w+1}-2^{w}$
$=-\frac{3}{2},$
with equality if $c=2^{w}-2$.
This completes the proof in all cases. ∎
3.4 Proof of Corollary 1.3
To prove (4), let $N\geq 2$ and choose $m\in\mathbb{N}$ such that $2^{m}\leq N\leq 2^{m+1}$. Since $q_{\infty}(Q_{N})$ is non-increasing in $N$, (2) gives
$N^{3/4}q_{\infty}(Q_{N})\leq(2^{m+1})^{3/4}q_{\infty}(Q_{2^{m}})\leq(2^{m+1})^{3/4}\cdot 2^{-3m/4-5/4}=2^{-1/2}.$
This proves (4) for general $N$.
If $N\geq 64$, then $m\geq 6$, and we can use (3) instead of (2); the same analysis then gives the improved constants for $q_{\infty}(Q_{N})$.
Finally, (5) follows immediately from (4) together with the general bound $h_{\infty}(Q_{N})\geq 1/(2\sqrt{N})$ as given in [3, Remark 2.4]. ∎

### v2

@@
 Keywords: Quasi-Monte Carlo, Sobol’ sequence, space-filling design, covering radius, separation radius, mesh ratio
-AMS subject classifications: 05B40, 11K36, 52C15
 1 Introduction
-Quasi-Monte Carlo (QMC) methods replace random sampling with carefully constructed deterministic point sets for numerical integration over the unit cube; see, e.g., [4, 5, 6, 13, 14, 19].
+Quasi-Monte Carlo (QMC) methods replace random sampling with carefully constructed deterministic point sets for numerical integration over the unit cube; see, e.g., [DKP22, DKS13, DP10, LePi14, N92, SJ94].
 While classical QMC theory focuses on discrepancy, which measures deviation from perfect equidistribution, it does not directly control local spacing between points.
 However, many applications require a more geometric notion of uniformity.
-In particular, tasks such as scattered data approximation [23], Gaussian process regression [22], kernel interpolation [18], and the design of computer experiments [7, 15, 17] benefit from point sets that ensure both well-controlled spacing and coverage.
+In particular, tasks such as scattered data approximation [W05], Gaussian process regression [T20], kernel interpolation [Sch95], and the design of computer experiments [FLS06, PM12, SWN03] benefit from point sets that ensure both well-controlled spacing and coverage.
 These properties are summarized by the term quasi-uniformity.
@@
 Intuitively, $h_{p}(Q)$ is the smallest radius such that closed $\ell^{p}$ balls around each point cover $0,1]^{d}$, whereas $q_{p}(Q)$ is the largest radius such that the corresponding open balls around the points do not overlap.
-As noted in [14, Section 6], [21], and [16], this geometric interpretation implies
+As noted in [N92, Section 6], [SS07], and [PZ23], this geometric interpretation implies
 $h_{p}(Q)\in\Omega(|Q|^{-1/d}),\qquad q_{p}(Q)\in O(|Q|^{-1/d}).$ (1)
@@
 Given that QMC point sets are natural candidates for generating uniform point sets, it is important to ask whether they are quasi-uniform.
-As noted in [24], however, the quasi-uniformity of classical QMC constructions had long been unresolved.
-This changed with Goda’s influential result [9], which shows that the two-dimensional Sobol’ sequence is not quasi-uniform.
-The Sobol’ sequence [20] is widely used in practice due to its efficient digital construction, extensibility to arbitrary sample sizes, and effectiveness in high-dimensional integration.
-It is available in standard software packages, such as Python’s QMCPy [1] and MATLAB’s Statistics and Machine Learning Toolbox.
-Sobol’ and Shukhman [21] conjectured that the separation radius of the first $N$ points of the $d$-dimensional Sobol’ sequence behaves like $N^{-1/d}$, which would imply quasi-uniformity.
-For $d=2$, Goda [9] showed that this conjecture fails for the sparse subsequence of the form $N=2^{2^{w}-1}$, and the case $N=2^{2^{w}}$ was subsequently analyzed in [3].
-Nevertheless, a complete description of the separation radius for general $N$ remains open, and addressing this gap is the focus of the present work.
+This question is non-trivial, since quasi-uniformity does not necessarily imply low discrepancy, nor does low discrepancy imply quasi-uniformity for $d\geq 2$ [DGLPSxx, Section 1.3].
+As noted in [WBG21], this issue remained open for a long time for classical QMC constructions.
+This situation changed with Goda’s influential result [G24a], which shows that the two-dimensional Sobol’ sequence is not quasi-uniform.
+The Sobol’ sequence [So67] is widely used in practice due to its efficient digital construction, extensibility to arbitrary sample sizes, and effectiveness in high-dimensional integration.
+It is available in standard software packages, such as Python’s QMCPy [QMCpy] and MATLAB’s Statistics and Machine Learning Toolbox.
+Sobol’ and Shukhman [SS07] conjectured that the separation radius of the first $N$ points of the $d$-dimensional Sobol’ sequence behaves like $N^{-1/d}$, which would imply quasi-uniformity.
+For $d=2$, Goda [G24a] showed that this conjecture fails for the sparse subsequence of the form $N=2^{2^{w}-1}$, and the case $N=2^{2^{w}}$ was subsequently analyzed in [DGSxx].
+Nevertheless, a complete characterization of the separation radius in the primary case $N=2^{m}$ remained open, and closing this gap is the aim of the present work.
 Main results. Let $Q_{N}$ denote the first $N$ points of the two-dimensional Sobol’ sequence (Definition 2.1). Our main contributions are as follows:
-  • We provide an exact formula for the $\ell^{\infty}$-separation radius of the dyadic prefixes $Q_{2^{m}}$.
+  • We provide an exact formula for the $\ell^{\infty}$-separation radius of $Q_{2^{m}}$ for all $m$.
   • We prove that $q_{\infty}(Q_{N})\in O(N^{-3/4})$ and $\rho_{\infty}(Q_{N})\in\Omega(N^{1/4})$.
@@
 $q_{\infty}(Q_{N})=2^{-m-1}.$
-Otherwise, decomposing $m$ as $m=2^{v}+2^{w}+c$ with integers $v>w\geq 0$ and $0\leq c<2^{w}$, then
+Otherwise, decomposing $m$ as $m=2^{v}+2^{w}+c$ with integers $v>w\geq 0$ and $0\leq c<2^{w}$ (note that the case $(w,c)=(v-1,2^{v-1}-1)$ is excluded since this falls into the first case as $m=2^{v+1}-1$), then
 $q_{\infty}(Q_{N})=2^{-{2^{v}}-2^{w}}.$
@@
 $q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-5/4},$ (2)
-$q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-3/2}\quad\text{for }m\neq 1,5,$ (3)
+$q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-3/2}\qquad\text{for }m\neq 1,5,$ (3)
 with equality in the second inequality if $m=2^{v}-2$ for some $v\geq 2$.
@@
 where $C_{1}=C_{2}=2^{-1/2}$. Moreover, for $N\geq 64$, the constants can be improved to $C_{1}=2^{-3/4}$ and $C_{2}=2^{-1/4}$.
+Sobol’ sequences are most commonly and effectively used when the number of points is a power of two ($N=2^{m}$).
+Therefore, our results address the primary use case and are sufficient for the vast majority of practical applications.
+Despite this practical sufficiency, the exact values of the separation radius for arbitrary $N$ remain a natural mathematical question, which we investigate in Section 4.
 Related work.
 In recent years, the quasi-uniformity of QMC point sets has been the subject of intensive investigation.
-There are two major classes of QMC constructions: digital nets and sequences [6, 14], and lattice point sets (including their infinite analogue, Kronecker sequences) [4, 19].
+There are two major classes of QMC constructions: digital nets and sequences [DP10, N92], and lattice point sets (including their infinite analogue, Kronecker sequences) [DKP22, SJ94].
 For lattice point sets, quasi-uniformity has been actively studied.
-In one dimension, the separation radius of Kronecker sequences is completely characterized in [8].
-More general results in higher dimensions are provided in [2], including bounded mesh ratios for two-dimensional Fibonacci lattices, existence results for $d$-dimensional lattice rules, and explicit constructions for $d$-dimensional Kronecker sequences.
+In one dimension, the separation radius of Kronecker sequences is completely characterized in [G24b].
+More general results in higher dimensions are provided in [DGLPSxx], including bounded mesh ratios for two-dimensional Fibonacci lattices, existence results for $d$-dimensional lattice rules, and explicit constructions for $d$-dimensional Kronecker sequences.
 Turning to digital nets and sequences, the one-dimensional van der Corput sequence in base $b$ is known to be quasi-uniform since its first $b^{m}$ points are $\{i/b^{m}\mid 0\leq i<b^{m}\}$.
-In higher dimensions, the covering radius, often referred to as dispersion, has been extensively studied [14, Chapter 6].
+In higher dimensions, the covering radius, often referred to as dispersion, has been extensively studied [N92, Chapter 6].
 In particular, for $(t,d)$-sequences in base bb—well-known examples include the Sobol’, Faure, and Niederreiter sequences—, the covering radius is known to attain the optimal order $\Theta(N^{-1/d})$ for any dimension $d$.
 Thus, the problem of establishing quasi-uniformity reduces to verifying whether the separation radius also scales as $N^{-1/d}$.
-For $d=2$, as stated, the Sobol’ sequence is not quasi-uniform [9].
-The separation radius of several two-dimensional digital nets was studied in [10, 11].
-Numerical experiments therein suggest that the Larcher–Pillichshammer nets [12] are quasi-uniform.
-This was theoretically proved by Dick, Goda, and Suzuki [3], who introduced an algebraic criterion for well-separated digital nets.
+For $d=2$, as stated, the Sobol’ sequence is not quasi-uniform [G24a].
+The separation radius of several two-dimensional digital nets was studied in [GHSK08, GK09].
+Numerical experiments therein suggest that the Larcher–Pillichshammer nets [LP01] are quasi-uniform.
+This was theoretically proved by Dick, Goda, and Suzuki [DGSxx], who introduced an algebraic criterion for well-separated digital nets.
 To our knowledge, this remains the only explicit construction of low-discrepancy and quasi-uniform digital nets for $d\geq 2$.
-The paper [3] also shows the non-optimality of the separation radius for some two-dimensional digital nets and Fibonacci polynomial lattices, as well as for $b$-dimensional Faure sequences in prime base $b$.
-From a different perspective, Pronzato and Zhigljavsky [16] constructed quasi-uniform infinite sequences via a greedy packing algorithm, ensuring a mesh ratio of at most $2$ for the first $N$ points, $N\geq 2$.
+The paper [DGSxx] also shows the non-optimality of the separation radius for some two-dimensional digital nets and Fibonacci polynomial lattices, as well as for $b$-dimensional Faure sequences in prime base $b$.
+From a different perspective, Pronzato and Zhigljavsky [PZ23] constructed quasi-uniform infinite sequences via a greedy packing algorithm, ensuring a mesh ratio of at most $2$ for the first $N$ points, $N\geq 2$.
 However, these sequences do not necessarily maintain low discrepancy in dimensions $d\geq 2$.
-Organization. Preliminaries and notation are collected in Section 2. Section 3 provides the necessary lemmas, the proof of Theorem 1.1, and the derivation of Corollaries 1.2 and 1.3.
+Organization.
+Preliminaries and notation are collected in Section 2.
+Section 3 provides the necessary lemmas, the proof of Theorem 1.1, and the derivation of Corollaries 1.2 and 1.3.
+Results for arbitrary NN are given in Section 4.
 2 Preliminaries
@@
 where $n_{1},\dots,n_{m}\in\{0,1\}$ are identified with elements of $\mathbb{F}_{2}$.
-For a vector $\boldsymbol{z}=(z_{1},\dots,z_{m})^{\top}\in\mathbb{F}_{2}^{m}$, we denote by $\boldsymbol{z}[i]\in\{0,1\}$ its $i$th component $z_{i}$, and for $i<j$ we define the slice $\boldsymbol{z}[i\!:\!j]=(z_{i},\dots,z_{j})^{\top}$.
+For a vector $\vec{z}=(z_{1},\dots,z_{m})^{\top}\in\mathbb{F}_{2}^{m}$, we denote by $\vec{z}[i]\in\{0,1\}$ its $i$th component $z_{i}$, and for $i<j$ we define the slice $\vec{z}[i\!:\!j]=(z_{i},\dots,z_{j})^{\top}$.
 Finally, we set
-$\phi(\boldsymbol{z})=\boldsymbol{z}[1]2^{-1}+\cdots+\boldsymbol{z}[m]2^{-m}.$
+$\phi(\vec{z})=\vec{z}[1]2^{-1}+\cdots+\vec{z}[m]2^{-m}.$
 For a matrix $P=(P_{ij})_{1\leq i,j\leq m}$, we use the notation $P[i][j]:=P_{ij}$ for convenience, and define
@@
 The subscript $m$ will be omitted whenever it does not cause confusion.
-Two-dimensional Sobol’ sequence. The two-dimensional Sobol’ sequence is defined as follows. For the definition in general dimension $d$, we refer the reader to [6, Chapter 8].
+Two-dimensional Sobol’ sequence. The two-dimensional Sobol’ sequence is defined as follows. For the definition in general dimension $d$, we refer the reader to [DP10, Chapter 8].
 Definition 2.1.
-Let $n\in\mathbb{N}_{0}$ and choose $m\in\mathbb{N}$ such that $n<2^{m}$.
 The two-dimensional Sobol’ sequence $(\boldsymbol{x}_{n})_{n\in\mathbb{N}}$ is a sequence of points in $0,1)^{2}$, where the $n$th point is given by
-$\boldsymbol{x}_{n}:=(\phi(\vec{n}),\phi(P_{m}\vec{n})).$
+$\boldsymbol{x}_{n}:=(\phi(\vec{n}),\phi(P_{m}\vec{n})),$
+where $m\in\mathbb{N}$ is chosen such that $n<2^{m}$.
 This definition does not depend on the choice of $m$: increasing $m$ simply pads $\vec{n}$ with leading zeros, and since $P_{m}$ is upper triangular, $\phi(\vec{n})$ and $\phi(P_{m}\vec{n})$ remain unchanged.
+Note that some implementations generate Sobol’ sequences in Gray code order rather than in the radical inverse order used here.
+This difference affects only the ordering of the points; for $N=2^{m}$, the resulting point sets are identical.
 This construction is generalized to the notions of digital nets and digital sequences.
-The uniformity of these point sets is usually measured by a quantity called the $t$-value (see [14, Chapter 4] or [6, Chapter 4]).
-It is known that the $t$-value of the two-dimensional Sobol’ sequence attains the best possible value, namely zero [6, Section 8.1].
+The uniformity of these point sets is usually measured by a quantity called the $t$-value (see [N92, Chapter 4] or [DP10, Chapter 4]).
+It is known that the $t$-value of the two-dimensional Sobol’ sequence attains the best possible value, namely zero [DP10, Section 8.1].
 This discussion can be formulated rigorously in the following proposition.
@@
     $(P_{V+W}\boldsymbol{1}_{V+W})[i]=1\iff i=W,V\text{ or }V+W$.
+  8. (viii)
+    $(P_{V-1}\boldsymbol{1}_{V-1})[i]=\boldsymbol{1}_{V-1}$.
 Proof.
@@
 $(P_{m}\boldsymbol{1}_{m})[i]\equiv\bigoplus_{j=1}^{m}\binom{j-1}{i-1}\equiv\binom{m}{i}\pmod{2}$.
-Using this fact, Items (vi) and (vii) also follow from Lucas’s theorem. ∎
+Using this fact, Items (vi)(viii) also follow from Lucas’s theorem. ∎
 The following results assert that the binary representations of two close points are related.
@@
 3.2 Proof of the main theorem
-  • The case $m=2^{v}-1$ is essentially treated in [9]. As shown in the proof of [9, Theorem 2.2], we have
-    $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}-1}\|_{\infty}=2^{-m}.$
+  • The case $m=2^{v}-1$ is essentially treated in [G24a]. As shown in the proof of [G24a, Theorem 2.2], we have
+    $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}-2}\|_{\infty}=2^{-m}.$
     On the other hand, by the construction of the Sobol’ sequence, $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}\geq 2^{-m}$ holds for any $p\neq q$.
@@
     $q_{\infty}(Q_{2^{m}})=2^{-m-1}.$
-  • The case $m=2^{v}$ is essentially treated in [3].
-From the proof of [3, Theorem 4.2] with $L$ being the identity matrix, we deduce $\boldsymbol{x}_{2^{m}+1}=(1/2+1/2^{m},1/2+1/2^{m})$, which implies $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}+1}\|_{\infty}=2^{-m}$.
+  • The case $m=2^{v}$ is essentially treated in [DGSxx].
+From the proof of [DGSxx, Theorem 4.2] with $L$ being the identity matrix, we deduce $\boldsymbol{x}_{2^{m-1}+1}=(1/2+1/2^{m},1/2-1/2^{m})$, which implies $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m-1}+1}\|_{\infty}=2^{-m}$.
 Hence, in the same manner as the previous case, we obtain
@@
 3.2.1 Proof of the upper bound
-Let $p=2^{V+W-1}+2^{W-1}$ and $q=2^{V+W}-2^{W}$. To prove $q_{\infty}(Q_{2^{m}})\leq 2^{-V-W}$, it suffices to show that $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-V-W+1}$.
-First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=W$ or $i=V+W$, it follows from Lemma 3.1 (i) and (ii) that
-$x_{p,1}$ $=\phi(\vec{p})=2^{-W}+2^{-V-W},$
-$x_{p,2}$ $=\phi(P_{m}\vec{p})=\sum_{i=1}^{V+W}(P[i][W]\oplus P[i][V+W])2^{-i}=\sum_{i=V+1}^{V+W}2^{-i}=2^{-V}-2^{-V-W}.$
-Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ for $W+1\leq i\leq V+W$, we have
-$P_{m}\vec{q}$ $=P_{m}(\mathrm{concat}(\boldsymbol{1}_{W},\boldsymbol{0}_{m-W})\oplus\mathrm{concat}(\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}))$
-$=\mathrm{concat}(P_{W}\boldsymbol{1}_{W},\boldsymbol{0}_{m-W})\oplus\mathrm{concat}(P_{V+W}\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}),$
+Let $p=2^{V+W-1}+2^{V-1}$ and $q=2^{V+W}-2^{V}$. To prove $q_{\infty}(Q_{2^{m}})\leq 2^{-V-W}$, it suffices to show that $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-V-W+1}$.
+First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=V$ or $i=V+W$, it follows from Lemma 3.1 (i) and (ii) that
+$x_{p,1}$ $=\phi(\vec{p})=2^{-V}+2^{-V-W},$
+$x_{p,2}$ $=\phi(P_{m}\vec{p})=\sum_{i=1}^{V+W}(P[i][V]\oplus P[i][V+W])2^{-i}=\sum_{i=W+1}^{V+W}2^{-i}=2^{-W}-2^{-V-W}.$
+Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ for $V+1\leq i\leq V+W$, we have
+$P_{m}\vec{q}$ $=P_{m}(\mathrm{concat}(\boldsymbol{1}_{V},\boldsymbol{0}_{m-V})\oplus\mathrm{concat}(\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}))$
+$=\mathrm{concat}(P_{V}\boldsymbol{1}_{V},\boldsymbol{0}_{m-V})\oplus\mathrm{concat}(P_{V+W}\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}),$
 where “concat” denotes vertical concatenation of column vectors. Thus, by Lemma 3.1 (vi) and (vii), we obtain
-$x_{q,1}$ $=\phi(\vec{q})=\sum_{i=W+1}^{V+W}2^{-i}=2^{-W}-2^{-V-W},$
-$x_{q,2}$ $=\phi(P_{m}\vec{q})=2^{-V}+2^{-V-W}.$
+$x_{q,1}$ $=\phi(\vec{q})=\sum_{i=V+1}^{V+W}2^{-i}=2^{-V}-2^{-V-W},$
+$x_{q,2}$ $=\phi(P_{m}\vec{q})=2^{-W}+2^{-V-W}.$
 Hence,
@@
 $(P\vec{\Delta})[V]=\vec{\Delta}[V]=0.$
-Then, Corollary 3.3 (i) implies
+Then, Corollary 3.3 (i) with $\ell=V+1$ implies
 $\vec{\Delta}[1\!:\!V]=(P\vec{\Delta})[1\!:\!V]=\boldsymbol{0}.$
@@
 Case 2: $\vec{\Delta}[V]=1$ and $\vec{\Delta}[V-1]=0$.
-In this case, Corollary 3.3 (ii), together with the assumption $\phi(\vec{p})<\phi(\vec{q})$, implies
+In this case, Corollary 3.3 (ii) with $k=V$ and $\ell=V+W$, together with the assumption $\phi(\vec{p})<\phi(\vec{q})$, implies
 $\vec{\Delta}[V\!:\!V+W-1]=\boldsymbol{1},$ (8)
 $\vec{\Delta}[V+W\!:\!m]\neq\boldsymbol{0},$ (9)
-$\vec{p}[V]=0,\quad\vec{q}[V]=1,$ (10)
+$\vec{p}[V]=0,\qquad\vec{q}[V]=1,$ (10)
 $\vec{p}[V+W]\geq\vec{q}[V+W].$ (11)
@@
 Case 2-1: $\vec{\Delta}[V+W]=0$.
-In this case, using (8) and Lemma 3.1 (v), we have
-$(P\vec{\Delta})[V+1\!:\!V+W-1]$ $=P[V+1\!:\!V+W-1][V+1\!:\!V+W-1]\cdot\vec{\Delta}[V+1\!:\!V+W-1]$
-$\quad\oplus P[V+1\!:\!V+W-1][V+W\!:\!V+W]\cdot\vec{\Delta}[V+W]$
-$\quad\oplus P[V+1\!:\!V+W-1][V+W+1\!:\!m]\cdot\vec{\Delta}[V+W+1\!:\!m]$
+In this case, using (12), (8), Lemma 3.1 (v), and Lemma 3.1 (viii), we have
+$\boldsymbol{1}$ $=(P\vec{\Delta})[V+1\!:\!V+W-1]$
+$=P[V+1\!:\!V+W-1][V+1\!:\!V+W-1]\cdot\vec{\Delta}[V+1\!:\!V+W-1]$
+$\qquad\oplus P[V+1\!:\!V+W-1][V+W\!:\!V+W]\cdot\vec{\Delta}[V+W]$
+$\qquad\oplus P[V+1\!:\!V+W-1][V+W+1\!:\!m]\cdot\vec{\Delta}[V+W+1\!:\!m]$
 $=P_{W-1}\boldsymbol{1}\oplus\boldsymbol{0}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m]$
-$=\boldsymbol{1}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m],$
-where $P^{\prime}:=P[V+1\!:\!V+W-1][V+W+1\!:\!m]$. Combined with (12), this gives
+$=\boldsymbol{1}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m].$
+where $P^{\prime}:=P[V+1\!:\!V+W-1][V+W+1\!:\!m]$. This gives
 $P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m]=\boldsymbol{0}.$ (14)
@@
 If $N\geq 64$, then $m\geq 6$, and we can use (3) instead of (2); the same analysis then gives the improved constants for $q_{\infty}(Q_{N})$.
-Finally, (5) follows immediately from (4) together with the general bound $h_{\infty}(Q_{N})\geq 1/(2\sqrt{N})$ as given in [3, Remark 2.4]. ∎
+Finally, (5) follows immediately from (4) together with the general bound $h_{\infty}(Q_{N})\geq 1/(2\sqrt{N})$ as given in [DGSxx, Remark 2.4]. ∎
+4 Separation radius for general NN
+Building on our previous result for the exact separation radius $q_{\infty}(Q_{2^{m}})$ for a power-of-two number of points, this section establishes an upper bound for $q_{\infty}(Q_{N})$ for an arbitrary number of points $N$.
+It is important to note that the analysis herein applies exclusively to the Sobol’ sequence in radical inverse order (Definition 2.1).
+To motivate our theoretical result, we first conducted a numerical experiment to identify the pairs of indices $(p,q)$ that successively update the minimum distance $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}$ for $q\leq 2^{15}$.
+Note that, since the Sobol’ sequence starts from index 0, the point set corresponding to $(p,q)$ consists of $N=q+1$ points.
+We recall that the separation radius is the half of the minimum distance.
+The results of this search, summarized in Table 1, motivate the following theorem, which provides an analytical result consistent with our numerical observations.
+Table 1: Indices $(p,q)$ for which the distance $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}$ successively achieves the minimum distance for the Sobol’ sequence, shown for $q\leq 2^{15}$.
+The fourth row indicates the corresponding case in Theorem 4.1 that describes the pair.
+The last two rows show the corresponding number of points $N=q+1$ and the separation radius $q_{\infty}(Q_{N})=\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}/2$, respectively.
+pp 0 11 22 11 $24$ $40$ 88 11 $504$ $640$ $2176$ $128$
+$q$ 11 22 44 99 $36$ $48$ $112$ $129$ $516$ $768$ $3840$ $32512$
+$\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}$ $^{-1} 22$$^{-2}$ $2^{-3}$ $2^{-4}$ $3\cdot 2^{-6}$ $2^{-5}$ $2^{-7}$ $2^{-8}$ $3\cdot 2^{-10}$ $2^{-9}$ $2^{-11}$ $2^{-15}$
+Thm. 4.1 (ii) (i) (iii) (iv) (ii) (i) (iii) (iv) (iv) (ii)
+NN 22 33 55 $10$ $37$ $49$ $113$ $130$ $517$ $769$ $3841$ $32513$
+$q_{\infty}(Q_{N})$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$ $3\cdot 2^{-7}$ $2^{-6}$ $2^{-8}$ $2^{-9}$ $3\cdot 2^{-11}$ $2^{-10}$ $2^{-12}$ $2^{-16}$
+Theorem 4.1.
+Let $Q=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots\}$ be the two-dimensional Sobol’ sequence and $Q_{N}$ be its first $N$ points. Let $v>w\geq 1$ be integers. Then we have the following.
+  1. (i)
+    Case where $m=2^{v}$: Let $p=1$, $q=2^{2^{v}-1}+1$, and $N=q+1$. Then
+    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-2^{v}-1}$
+  2. (ii)
+    Case where $m=2^{v}-1$: Let $p=2^{2^{v-1}-1}$, $q=2^{2^{v}-1}-2^{2^{v-1}}$, and $N=q+1$. Then
+    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-2^{v}}.$
+  3. (iii)
+    Case where $m=2^{v}+2$: Let $p=2^{2^{v}+1}-8$, $q=2^{2^{v}+1}+4$, and $N=q+1$. Then
+    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=3\cdot 2^{-2^{v}-3}$
+  4. (iv)
+    Case where $m=2^{v}+2^{w}$: Let $p=2^{2^{v}+2^{w}-1}+2^{2^{v}-1}$, $q=2^{2^{v}+2^{w}}-2^{2^{v}}$, and $N=q+1$. Then
+    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-2^{v}-2^{w}}.$
+Proof.
+In this proof, we use the notation $V=2^{v}$ and $W=2^{w}$ as in Section 3.
+(i): This is shown in the second case of Section 3.2.
+(ii): Let $V^{\prime}=2^{v-1}$. First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=V^{\prime}$ , it follows from Lemma 3.1 (i) that
+$x_{p,1}$ $=\phi(\vec{p})=2^{-V^{\prime}},$
+$x_{p,2}$ $=\phi(P_{V-1}\vec{p})=\sum_{i=1}^{V^{\prime}}P[i][V^{\prime}]2^{-i}=\sum_{i=1}^{V^{\prime}}2^{-i}=1-2^{-V^{\prime}}.$
+Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. In what follows, “concat” denotes vertical concatenation of column vectors. Since $\vec{q}[i]=1$ for $V^{\prime}+1\leq i\leq V-1$, we have
+$P_{V-1}\vec{q}$ $=P_{V-1}(\boldsymbol{1}_{V-1}\oplus\mathrm{concat}(\boldsymbol{1}_{V^{\prime}},\boldsymbol{0}_{V-V^{\prime}-1}))$
+$=P_{V-1}\boldsymbol{1}_{V-1}\oplus\mathrm{concat}(P_{V^{\prime}}\boldsymbol{1}_{V^{\prime}},\boldsymbol{0}_{V-V^{\prime}-1}))$
+[…]
+$=\mathrm{concat}(\boldsymbol{1}_{V^{\prime}-1},\boldsymbol{0}_{1},\boldsymbol{1}_{V-V^{\prime}-1}),$
+where the third equality follows from Lemma 3.1 (vi) and (viii). Thus we obtain
+$x_{q,1}$ $=\phi(\vec{q})=\sum_{i=V^{\prime}+1}^{V-1}2^{-i}=2^{-V^{\prime}}-2^{-V+1}$
+$x_{q,2}$ $=\phi(P_{V-1}\vec{q})=\sum_{i=1}^{V-1}2^{-i}-2^{-V^{\prime}}=1-2^{-V^{\prime}}-2^{-V+1}.$
+Hence $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|=2^{-V+1}$.
+(iii): First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. We have $\vec{p}[i]=1$ for $4\leq i\leq V+1$. From Lemma 3.1 (vii) with $(V,W)=(2,1)$ and $(V,1)$, we have
+$(P_{3}\boldsymbol{1}_{3})[i]$ $=1\iff i=1,2,3$
+$(P_{V+1}\boldsymbol{1}_{V+1})[i]$ $=1\iff i=1,V,V+1.$
+Since we have assumed that $v\geq 2$ and thus $V\geq 4$, we have
+$(P_{V+2}\vec{p})[i]=1\iff i=2,3,V,V+1,$
+and thus
+$x_{p,1}$ $=\phi(\vec{p})=\sum_{i=4}^{V+1}2^{-i}=2^{-3}-2^{-V-1},$
+$x_{p,2}$ $=\phi(P_{V+2}\vec{p})=2^{-2}+2^{-3}+2^{-V}+2^{-V-1}.$
+Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ if and only if $i=3$ or $i=V+2$, it follows from Lemma 3.1 (ii) with $(V,W)=(2,1)$ and $(V,2)$ that
+$x_{q,1}$ $=\phi(\vec{q})=2^{-3}+2^{-V-2}$
+$x_{q,2}$ $=\phi(P_{V+2}\vec{q})=\sum_{i=1}^{V+2}(P[i][3]\oplus P[i][V+2])2^{-i}=2^{-2}+2^{-3}+2^{-V-1}+2^{-V-2}.$
+Hence $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|=3\cdot 2^{-V-2}$.
+(iv): This is shown in Section 3.2.1. ∎
+Establishing a lower bound on the separation radius or minimum distance for general $N$ would necessitate a significantly more complex argument than the one used for the proof in Section 3.2.2.
+We therefore propose the conjecture that, for $q\geq 4$, a pair $(\boldsymbol{x}_{p},\boldsymbol{x}_{q})$ updates the minimum distance if and only if it is one of the candidates specified in Theorem 4.1.
+5 Conclusion
+We established a complete characterization of the separation radius for the case $N=2^{m}$ and investigated its numerical behavior for general $N$.
+Our analysis reveals that the two-dimensional Sobol’ sequence has a suboptimal mesh ratio that grows at least as $N^{1/4}$.
+While digital scrambling may alleviate the poor growth of the separation radius, examining this possibility is left for future work.
+In sharp contrast to the two-dimensional case (Theorem 1.1), Sobol’ points in dimensions $d\geq 3$ did not exhibit a clear systematic pattern in their separation radius.
+For $d=3$, in particular, the observed rate lies between $N^{-1/3}$ and $N^{-1/2}$ up to $N\leq 2^{15}$.
+Understanding the behavior in higher dimensions and establishing dimension-dependent bounds remain important directions for future research.
+Acknowledgments
+The author acknowledge the use of OpenAI’s ChatGPT 5 and Google’s Gemini 2.5 for editing and language polishing of the manuscript.